OpenStudy (anonymous):

Which does not show a direct variation between x and y? A. y = B. y = 2x C. y = 0.5x D. y =

OpenStudy (anonymous):

D. y= x/y

OpenStudy (anonymous):

what is in A?

OpenStudy (mathmate):

Direct variation is in the form y=kx where k is a real numeric constant. Examples: y=4x, y=sqrt(23)x y=(3/2)x Examples that are NOT direct variations y=(4/3)x*x y=x/y^2 y=x+3 y=2x-pi

OpenStudy (anonymous):

\[y=\frac{ k }{ x }\] here x and y are in indirect variation.